# Restless Bandit Problem with Rewards Generated by a Linear Gaussian Dynamical System

### v1

@@ Abstract
The stochastic multi-armed bandit problem studies decision-making under uncertainty. In the problem, the learner interacts with an environment by choosing an action at each round, where a round is an instance of an interaction. In response, the environment reveals a reward, which is sampled from a stochastic process, to the learner. The goal of the learner is to maximize cumulative reward. A specific variation of the stochastic multi-armed bandit problem is the restless bandit, where the reward for each action is sampled from a Markov chain. The restless bandit with a discrete state-space is a well-studied problem, but to the best of our knowledge, not many results exist for the continuous state-space version which has many applications such as hyperparameter optimization. In this work, we tackle the restless bandit with continuous state-space by assuming the rewards are the inner product of an action vector and a state vector generated by a linear Gaussian dynamical system. To predict the reward for each action, we propose a method that takes a linear combination of previously observed rewards for predicting each action’s next reward. We show that, regardless of the sequence of previous actions chosen, the reward sampled for any previously chosen action can be used for predicting another action’s future reward, i.e. the reward sampled for action 1 at round $t-1$ can be used for predicting the reward for action $2$ at round $t$. This is accomplished by designing a modified Kalman filter with a matrix representation that can be learned for reward prediction. Numerical evaluations are carried out on a set of linear Gaussian dynamical systems.
@@ 1 Introduction
A more specific variation of the non-stationary SMAB are rewards are generated by $s$-step autoregressive models, i.e. an action’s sampled reward $X_{t}$ is a linear combination of rewards $X_{t-s},\dots,X_{t-1}$ where $s$ is the autoregressive model order. Two key results that have tackled this SMAB environment are [2], [3], and [4]. [2] studied the performance of a number of algorithms for rewards generated by Brownian motion. In [3], the authors consider when the rewards for each action is generated by a known 1-step autoregressive process. In [4], they address SMAB environments modeled as a unknown 1-step autoregressive or a known $s$-step autoregressive. A key application of autoregressive models is presented in [5], where the work tunes the hyperparameters, such as the gradient descent’s learning rate, during the training process of a reinforcement learning neural network. Finally, another perspective to $s$-step autoregressive models is [6] where rewards $X_{t}$ and a context $\theta_{t}$ are generated by a Linear Gaussian Dynamical System (LGDS). The authors prove that a linear combination of previously observed contexts $\theta_{t-s},\dots,\theta_{t-1}$ can be used to predict the reward $X_{t}$, a perspective similar to the environments considered in [3] and [4].
Figure 1: In this work, we illustrate that for the proposed LGDS, the linear combination of any sequence of observed rewards can be used to predict each action’s reward. For example, if we take the previously observed rewards $X_{t-3}$, $X_{t-2}$, $X_{t-1}$ based on the chosen actions $A_{t-3}=3$, $A_{t-2}=2$, $A_{t-1}=1$, we can predict the reward for actions $a=1,2,3$ using the linear combination of previously observed rewards. The linear combination is based on an estimated vector $\hat{G}_{a|\mathbf{c}}$, $a\in\{1,2,3\}$.
Our work proposes a discrete-time restless bandit with continuous state-space by assuming the state and rewards are generated by a LGDS. This paper extends the results in [6] where now the context is absent. The contributions of our paper are as follows.
@@
  • We introduce a discrete-time, continuous state-space, restless bandit where the rewards are generated by a LGDS in Section 2.
  • We prove that we can predict the reward for each action by using a linear combination of observed rewards. The coefficients for the linear combination are from the identified modified Kalman filter matrix representation. (See Section 3)
  • We provide a proof of the error bound of the reward prediction for the identified modified Kalman filter. The idea is inspired by [7] for identifying the Kalman filter, where now we assume that the measurements of the LGDS, a linear combination of the system’s state, can change each round.
  • Using the proved error bound of the reward prediction, we propose the algorithm Optimistic Sensor Selector (OSS). The algorithm chooses the action that maximizes the sum of the reward prediction and its error. (See Section 4)
  • For numerical results in Section 5, we apply OSS to a parameterized LGDS to illustrate its numerical performance. Here, we compare OSS to Upper Confidence Bound (UCB) algorithm [8] and Sliding Window UCB (SW-UCB) [9] algorithm, two well-known SMAB algorithms, and for which LGDS OSS performs best.
@@
An environment considered in non-stationary SMAB is the restless bandit in [10]. In the restless bandit, the reward for each action is the function of a state that is generated by a Markov chain. Whenever the learner chooses an action, the learner observes a Markov chain’s state and a reward. This paper focuses on the case when the transition matrix of the Markov chain is unknown. Previous results in the discrete state-space Markov chain that use an approach similar to UCB are [11, 12, 13, 14, 15]. [16] uses Thompson sampling, i.e. sampling parameters based on a priori distribution of Markov chain, for action selection. We avoid comparisons with these previous results since the states of the Markov chain are discrete, whereas the results presented in this paper focus on when the states are continuous. This allows us to tackle a different set of application, such as hyperparameter optimization for reinforcement learning neural networks, e.g. [5].
@@ 2 Problem Formulation
For this paper, the learner will interact with an environment modeled as a LGDS. We will consider the following LGDS where the rewards are the inner product of an action vector and the LGDS’s state:
@@
where $z_{t}\in\mathbb{R}^{d}$ is the state of the system and $X_{t}\in\mathbb{R}$ is the reward. The process noise $\xi_{t}\in\mathbb{R}^{d}$ and measurement noise $\eta_{t}\in\mathbb{R}$ are independent normally distributed random variables, i.e. $\xi_{t}\sim\mathcal{N}(0,Q)$ and $\eta_{t}\sim\mathcal{N}(0,\sigma)$. The vector $c_{A_{t}}\in\mathcal{A}=\{c_{a}\in\mathbb{R}^{d\times 1}\mid\left\|c_{a}\right\|_{2}\leq B_{c},a\in[k]\}$ where $B_{c}$ is known and $a\in[k]\triangleq\{1,2,\dots,k\}$ is the indexed action. Using similar notation as [17], actions that are chosen at round $t$ are denoted as $c_{A_{t}}\in\mathcal{A}$; otherwise, the action is denoted as $c_{a}\in\mathcal{A}$. We make the following assumptions on system (1).
@@ 3 Predicting the Reward of the LGDS
This section reviews the optimal $1$-step predictor of the rewards, in the mean-squared error sense, generated by LGDS (1): the Kalman filter. According to Assumption 2, the matrices of the LGDS (1) are unknown, implying that the Kalman filter needs to be identified. However, direct identification of the Kalman filter for the LGDS (1) is intractable. Therefore, we propose a modified Kalman filter to identify. Imposing the assumptions posed in the previous section, we prove that prediction error of the modified Kalman filter is lower than the variance of the reward $X_{t}$, making it a sufficiently useful predictor of the reward. The added benefit of the modified Kalman filter is that it is tractable to identify.
@@
The initial state $z_{0}\in\mathbb{R}^{d}$ of the LGDS (1) is sampled from a normal distribution with mean $\hat{z}_{0}=\lim_{\tau\rightarrow\infty}\Gamma^{\tau}\hat{z}$, where $\hat{z}\in\mathbb{R}^{d}$ is an arbitrary vector, and covariance matrix $P_{0}$. We assume for $P_{0}$ that $P_{0}=P_{\overline{a}}$, where $P_{\overline{a}}$ is the steady-state error covariance matrix, $P_{\overline{a}}=g\left(P_{\overline{a}},c_{\overline{a}}\right)$, $c_{\overline{a}}\in\mathcal{A}$.
@@ 3.1 Learning the Modified Kalman filter
Using Theorem 1 and inspired by the results presented in [7], we will learn the modified Kalman filter. Let parameter $s>0$ denote how far in the past the learner will look. We define the tuple $\mathbf{c}\triangleq\begin{pmatrix}c_{A_{t-s}}&\dots&c_{A_{t-1}}\end{pmatrix}$ as the sequence of actions chosen by the learner from rounds $t-s$ to $t-1$. The reward $X_{t}=\left\langle c_{a},z_{t}\right\rangle+\eta_{t}$ for action $a\in[k]$ can be expressed as a linear combination of rewards $X_{t-s},\dots,X_{t-1}$ generated by the tuple $\mathbf{c}$ using the matrices defined in the modified Kalman filter (5):
@@
where $Z_{t}$ (which has the iteration $Z_{t+1}=\Gamma Z_{t}\Gamma^{\top}+Q$) is the covariance of the LGDS’s (1) state $z_{t}$.
@@
4 Optimistic Sensor Selector LGDS Restless Bandit Problem
The section above provided a predictor, the modified Kalman filter, for the rewards generated by the LGDS (1). It also provided a methodology for identifying the predictor. Now that the reward can be predicted using an identified modified Kalman filter, we discuss how to use the predictor in Algorithm 1, Optimistic Sensor Selector (OSS). For each round $t$ in OSS, the learner will choose actions based on the following optimization problem
$\underset{a\in[k]}{\arg\max}~{}\hat{G}_{c_{a}\mid\mathbf{c}}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}\Xi_{t}\left(\mathbf{c}\right)+\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)\sqrt{\Xi_{t}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{t}\left(\mathbf{c}\right)},$ (12)
where with a probability of at least $(1-\delta)^{2}$, the following inequality is satisfied (see Theorem 4 in Appendix C.2):
$\hat{G}_{c_{a}\mid\mathbf{c}}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}\Xi_{t}\left(\mathbf{c}\right)-G_{c_{a}\mid\mathbf{c}}^{\top}\Xi_{t}\left(\mathbf{c}\right)\leq\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)\sqrt{\Xi_{t}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{t}\left(\mathbf{c}\right)}.$ (13)
The terms $e_{c_{a}\mid\mathbf{c}}\left(\delta\right)$ and $b_{c_{a}\mid\mathbf{c}}\left(\delta\right)$ are defined as
$e_{c_{a}\mid\mathbf{c}}\left(\delta\right)$ $\triangleq\sqrt{2B_{R}^{2}\log\left(\frac{1}{\delta}\frac{\det(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right))^{1/2}}{\det(\lambda I)^{1/2}}\right)}$ (14)
$b_{c_{a}\mid\mathbf{c}}\left(\delta\right)$ $\triangleq\sqrt{N_{a}}\frac{B_{c}B_{R}}{\delta}\sqrt{\mbox{tr}\left(I-\lambda V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\right)}+\lambda\sqrt{\mbox{tr}\left(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\right)}B_{G}.$ (15)
@@ 4 Optimistic Sensor Selector LGDS Restless Bandit Problem
  • Exploration term: $\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)\sqrt{\Xi_{t}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{t}\left(\mathbf{c}\right)}$
Parameter $\delta$ is a failure rate of the bound in (13) where $\delta$ values closer to 0 computes a larger bound (13). Parameter $s$ is the number of previously observed rewards $X_{t-\tau}$ ($\tau=1,\dots,s$) used for predicting the next reward. The number of models to learn increases exponentially as $s$ increases. Finally, $\lambda$ is a regularization parameter to ensure that (11) is invertible.
Input: $\delta\in(0,1)$, $\lambda>0$, $s\in\mathbb{N}$, $B_{c},B_{R},B_{G}\in\mathbb{R}^{+}$
@@
             $\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right),b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)\leftarrow 1/\epsilon$ where $\epsilon$ small
@@
             $A_{t}\leftarrow\underset{a\in\{1,2,\dots,k\}}{\arg\max}\hat{G}_{c_{a}\mid\mathbf{c}}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}\Xi_{t}\left(\mathbf{c}\right)+\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)\sqrt{\Xi_{t}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{t}\left(\mathbf{c}\right)}$
@@
Algorithm 1 Optimistic Sensor Selector (OSS)

### v2

@@ Abstract
Decision-making under uncertainty is a fundamental problem encountered frequently and can be formulated as a stochastic multi-armed bandit problem. In the problem, the learner interacts with an environment by choosing an action at each round, where a round is an instance of an interaction. In response, the environment reveals a reward, which is sampled from a stochastic process, to the learner. The goal of the learner is to maximize cumulative reward. In this work, we assume that the rewards are the inner product of an action vector and a state vector generated by a linear Gaussian dynamical system. To predict the reward for each action, we propose a method that takes a linear combination of previously observed rewards for predicting each action’s next reward. We show that, regardless of the sequence of previous actions chosen, the reward sampled for any previously chosen action can be used for predicting another action’s future reward, i.e. the reward sampled for action 1 at round $t-1$ can be used for predicting the reward for action $2$ at round $t$. This is accomplished by designing a modified Kalman filter with a matrix representation that can be learned for reward prediction. Numerical evaluations are carried out on a set of linear Gaussian dynamical systems and are compared with 2 other well-known stochastic multi-armed bandit algorithms.
@@ 1 Introduction
A more specific variation of the non-stationary SMAB are environments where the rewards are generated by $s$-step autoregressive models, i.e. an action’s sampled reward $X_{t}$ is a linear combination of rewards $X_{t-s},\dots,X_{t-1}$ where $s$ is the autoregressive model order. Two key results that have tackled this SMAB environment are [2, 3, 4]. [2] studied the performance of a number of algorithms for rewards generated by Brownian motion. In [3], the authors consider when the rewards for each action is generated by a known 1-step autoregressive process. In [4], they address SMAB environments modeled as an unknown 1-step autoregressive or a known $s$-step autoregressive. A key application of autoregressive models is presented in [5], where the work tunes the hyperparameters, such as the gradient descent’s learning rate, during the training process of a reinforcement learning based on neural networks. Finally, another perspective to $s$-step autoregressive models is [6] where the reward $X_{t}$ and a context $\theta_{t}$ are generated by a Linear Gaussian Dynamical System (LGDS), where a context is a partial observation of the LGDS’s state variables. The authors prove that a linear combination of previously observed contexts $\theta_{t-s},\dots,\theta_{t-1}$ can be used to predict the reward $X_{t}$, a perspective similar to the environments considered in [3] and [4].
Our work proposes a discrete-time restless bandit with continuous state-space by assuming the state and rewards are generated by a LGDS. This paper extends the results in [6] where now the context is no longer observed. The contributions of our paper are as follows.
@@
  • We introduce a SMAB environment where the rewards are generated by a LGDS in Section 2.
  • We prove that we can predict the reward for each action by using a linear combination of observed rewards. For example, for an environment with 3 actions, if a learner chose action $A_{t-2}=2$ at round $t-2$ and $A_{t-1}=1$ at round $t-1$, the learner can take a linear combination of the sampled rewards $X_{t-2}$ and $X_{t-1}$ to predict the reward for action $a=3$ at round $t$. The coefficients for the linear combination are from the identified modified Kalman filter matrix representation. We provide a proof of the error bound of the reward prediction for the identified modified Kalman filter. The idea is inspired by [7] for identifying the Kalman filter, where now we assume that the measurements of the LGDS, a linear combination of the system’s state variables, can change each round. (See Section 3)
  • Using the proved error bound of the reward prediction, we propose the algorithm Uncertainty-Based System Search (UBSS). The algorithm chooses the action that maximizes the sum of the reward prediction and its error. (See Section 4)
  • For numerical results in Section 5, we apply UBSS to a parameterized LGDS to illustrate its numerical performance. Here, we compare UBSS to Upper Confidence Bound (UCB) algorithm [8] and Sliding Window UCB (SW-UCB) [9] algorithm, two well-known SMAB algorithms, and for which LGDS UBSS performs best.
@@
One example of the non-stationary SMAB is the restless bandit where the reward for each action is the function of a state that is generated by a Markov chain [10]. Whenever the learner chooses an action, the learner observes a Markov chain’s state and a reward. This paper focuses on the case when the transition matrix of the Markov chain is unknown. Previous results in the discrete state-space Markov chain that use an approach similar to UCB are [11, 12, 13, 14, 15]. [16] uses Thompson sampling, i.e. sampling parameters based on a priori distribution of Markov chain, for action selection. We avoid comparisons with these previous results since the states of the Markov chain are discrete, whereas the results presented in this paper focus on when the states are continuous. This allows us to tackle a different set of application, such as hyperparameter optimization for reinforcement learning based on neural networks, e.g. [5].
@@ 2 Problem Formulation
The learner will interact with an environment modeled as a LGDS. We will consider the following LGDS:
@@
where the reward $X_{t}\in\mathbb{R}$ is the inner product of an action vector $c_{A_{t}}\in\mathcal{A}$ and the state $z_{t}\in\mathbb{R}^{d}$. The process noise $\xi_{t}\in\mathbb{R}^{d}$ and measurement noise $\eta_{t}\in\mathbb{R}$ are independent normally distributed random variables, i.e. $\xi_{t}\sim\mathcal{N}(0,Q)$ and $\eta_{t}\sim\mathcal{N}(0,\sigma^{2})$. The action vector $c_{A_{t}}\in\mathcal{A}=\{c_{a}\in\mathbb{R}^{d\times 1}\mid\left\|c_{a}\right\|_{2}\leq B_{c},a\in[k]\}$ where $B_{c}$ is known and $a\in[k]\triangleq\{1,2,\dots,k\}$ is the indexed action. Using similar notation as [17], actions that are realized at round $t$ are denoted as $c_{A_{t}}\in\mathcal{A}$ and unrealized actions are denoted as $c_{a}\in\mathcal{A}$. We make the following assumptions on system (1).
@@ 3 Predicting the Reward of the LGDS
This section reviews the optimal $1$-step predictor of the rewards, in the mean-squared error sense, generated by LGDS (1): the Kalman filter. According to Assumption 2, the matrices of the LGDS (1) are unknown, implying that the Kalman filter needs to be identified. However, to the best of our knowledge, no current results exist for direct identification of the Kalman filter when the LGDS’s (1) action vector $c_{A_{t}}\in\mathcal{A}$ can change each round. Therefore, we propose a modified Kalman filter to identify. Imposing the assumptions posed in the previous section, we prove that prediction error of the modified Kalman filter is lower than or equal to the variance of the reward $X_{t}$, making it possible to extract a signal to predict the reward for each action. The added benefit of the modified Kalman filter is that it is tractable to identify.
@@
The initial state $z_{0}\in\mathbb{R}^{d}$ of the LGDS (1) is sampled from a normal distribution with a mean $\hat{z}_{0}\in\mathbb{R}^{d}$ that is a solution of $\hat{z}_{0}=\Gamma\hat{z}_{0}$ and covariance matrix $P_{0}\in\mathbb{R}^{d\times d}$. We assume that $P_{0}=P_{\overline{a}}$, where $P_{\overline{a}}$ is the steady-state error covariance matrix, $P_{\overline{a}}=g\left(P_{\overline{a}},c_{\overline{a}}\right)$, $c_{\overline{a}}\in\mathcal{A}$. This assumption implies that the LGDS (1) is in a steady-state distribution.
@@ 3.1 Learning the Modified Kalman filter
Using Theorem 1 and inspired by the results presented in [7], we will learn the modified Kalman filter since the matrices and vectors in the LGDS (1) and its modified Kalman filter (5) are unknown. Let parameter $s>0$ denote how far in the past the learner will look. We define the tuple $\mathbf{c}\triangleq\begin{pmatrix}c_{A_{t-s}}&\dots&c_{A_{t-1}}\end{pmatrix}$ as the sequence of actions chosen by the learner from rounds $t-s$ to $t-1$. The reward $X_{t}=\left\langle c_{a},z_{t}\right\rangle+\eta_{t}$ for action $a\in[k]$ can be expressed as a linear combination of rewards $X_{t-s},\dots,X_{t-1}$ generated by the tuple $\mathbf{c}$ using the matrices defined in the modified Kalman filter (5):
@@
where $Z_{t}$ (which has the iteration $Z_{t+1}=\Gamma Z_{t}\Gamma^{\top}+Q$) is the covariance of the LGDS’s (1) state $z_{t}$. Results in the area of non-stationary SMAB have made similar assumptions (see [4]).
@@
4 Uncertainty-Based System Search Restless Bandit Problem
The section above provided a predictor, the modified Kalman filter, for the rewards generated by the LGDS (1). It also provided a methodology for identifying the predictor. Now that the reward can be predicted using an identified modified Kalman filter, we discuss how to use the predictor in Algorithm 1, Uncertainty-Based System Search (UBSS). The general scheme for UBSS is to 1) identify the predictor $G_{c_{a}\mid\mathbf{c}}$ for each action $c_{a}\in\mathcal{A}$ and 2) select actions that balances what the learner predicts will return the highest reward versus which actions the learner is the most uncertain due to the error of the predictor $\hat{G}_{c_{a}\mid\mathbf{c}}$. Therefore, for each round $t$ in UBSS, the learner will choose actions based on the following optimization problem
$\underset{a\in[k]}{\arg\max}~{}\hat{G}_{c_{a}\mid\mathbf{c}}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}\Xi_{t}\left(\mathbf{c}\right)+\left(e_{c_{a}\mid\mathbf{c}}\left(\delta_{e}\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta_{b}\right)\right)\sqrt{\Xi_{t}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{t}\left(\mathbf{c}\right)},$ (12)
where with a probability of at least $(1-\delta_{e})(1-\delta_{b})$, the following inequality is satisfied (see Theorem 9 in Appendix C.2):
$\hat{G}_{c_{a}\mid\mathbf{c}}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}\Xi_{t}\left(\mathbf{c}\right)-G_{c_{a}\mid\mathbf{c}}^{\top}\Xi_{t}\left(\mathbf{c}\right)\leq\left(e_{c_{a}\mid\mathbf{c}}\left(\delta_{e}\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta_{b}\right)\right)\sqrt{\Xi_{t}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{t}\left(\mathbf{c}\right)}.$ (13)
The terms $e_{c_{a}\mid\mathbf{c}}\left(\delta_{e}\right)$ and $b_{c_{a}\mid\mathbf{c}}\left(\delta_{b}\right)$ are defined as
$e_{c_{a}\mid\mathbf{c}}\left(\delta_{e}\right)$ $\triangleq\sqrt{2B_{R}^{2}\log\left(\frac{1}{\delta_{e}}\frac{\det(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right))^{1/2}}{\det(\lambda I)^{1/2}}\right)}$ (14)
$b_{c_{a}\mid\mathbf{c}}\left(\delta_{b}\right)$ $\triangleq\sqrt{N_{a}}\frac{B_{c}B_{R}}{\delta_{b}}\sqrt{\mbox{tr}\left(I-\lambda V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\right)}+\lambda\sqrt{\mbox{tr}\left(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\right)}B_{G}.$ (15)
@@ 4 Uncertainty-Based System Search Restless Bandit Problem
  • Exploration term: $\left(e_{c_{a}\mid\mathbf{c}}\left(\delta_{e}\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta_{b}\right)\right)\sqrt{\Xi_{t}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{t}\left(\mathbf{c}\right)}$
Parameters $\left(\delta_{e},\delta_{b}\right)$ are failure rates of the bound in (13) where $\left(\delta_{e},\delta_{b}\right)$ values closer to 0 computes a larger bound (13). Parameter $s$ is the number of previously observed rewards $X_{t-\tau}$ ($\tau=1,\dots,s$) used for predicting the next reward. The number of models to learn increases exponentially as $s$ increases. Finally, $\lambda$ is a regularization parameter to ensure that (11) is invertible.
Input: $\delta_{e},\delta_{b}\in(0,1)$, $\lambda>0$, $s\in\mathbb{N}$, $B_{c},B_{R},B_{G}\in\mathbb{R}^{+}$
@@
             $\left(e_{c_{a}\mid\mathbf{c}}\left(\delta_{e}\right),b_{c_{a}\mid\mathbf{c}}\left(\delta_{b}\right)\right)\leftarrow 1/\epsilon$ where $\epsilon$ small
@@
             $A_{t}\leftarrow\underset{a\in\{1,2,\dots,k\}}{\arg\max}\hat{G}_{c_{a}\mid\mathbf{c}}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}\Xi_{t}\left(\mathbf{c}\right)+\left(e_{c_{a}\mid\mathbf{c}}\left(\delta_{e}\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta_{b}\right)\right)\sqrt{\Xi_{t}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{t}\left(\mathbf{c}\right)}$
@@
Algorithm 1 Uncertainty-Based System Search (UBSS)